OpenStudy (anonymous):

h(x)=x-1 ------ x^3-9x find the domain and function

OpenStudy (welshfella):

the domain will not be all real numbers because there are values of x which will make the denominator = 0 which makes the value of the function indeterminate. You can find these values by solving x^3 - 9x = 0

OpenStudy (welshfella):

the domain will be all Real numbers except the 2 zeroes of the above equation.

OpenStudy (welshfella):

can you solve this equation?

OpenStudy (anonymous):

i was doing my test review the answer they gaive is (-inf,-3)U(-3,0)U(0,3)U(3,INF)

OpenStudy (welshfella):

yes thats the result in interval form x^3 - 9x = 0 x(x^2 - 9) = 0 x = 0 , 3 , -3 - there are 3 values not 2 ( i missed out the -3)

OpenStudy (anonymous):

NOW I SEE HOW THE ANSWER IS MADE

OpenStudy (welshfella):

so we start with minus infinity up to -3 - all these are in the domain except -3

OpenStudy (welshfella):

then thers the numbers between -3 and 0 included but not 0 - then numbers between 0 and 3 but not 0 and 3 , finally you have all numbers abobe ove 3 to infinity. The U means ;union'.

OpenStudy (anonymous):

THANKS I ALSO WILL POST ANOTHER QUESTION I WISH TO KNOW HOW TO GET THE ANSWER

OpenStudy (welshfella):

ok

OpenStudy (welshfella):

the parentheses ( and ) mean 'between but not including the first and last numbers'